OpenStudy (anonymous):

Which number sentence correctly solves this problem? Jimmy gave $550 to his favorite charity. He had already given $492 to this charity. How much money did Jimmy give to this charity? A. 550 + 492 = 942 B. 550 + 492 = 1,042 C. 550 - 492 = 58 D. 550 - 492 = 168

OpenStudy (mathmale):

Are you adding or subtracting here? Why? Which answers could not possibly be correct?

OpenStudy (anonymous):

you should know math-male

OpenStudy (imsleeziboii):

You are adding in this situation asiyah, therefore the equation is the two numbers being added together

OpenStudy (anonymous):

thankyou

OpenStudy (imsleeziboii):

you're welcome

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

ITS BETWEEN A AND B ASIYAH

OpenStudy (anonymous):

IT IS B

OpenStudy (imsleeziboii):

B**

OpenStudy (imsleeziboii):

the answer is B

OpenStudy (anonymous):

yeah learn how to add

OpenStudy (anonymous):



OpenStudy (anonymous):

Yup B

OpenStudy (anonymous):

550 + 492 = 1,042

OpenStudy (anonymous):

\[Which number sentence correctly solves this problem? Jimmy gave $550 \to his favorite charity. He had already given $492 \to this charity. How much money did Jimmy give \to this charity? A. 550 + 492 = 942 B. 550 + 492 = 1,042 C. 550 - 492 = 58 D.\]

OpenStudy (imsleeziboii):

wtf

OpenStudy (imsleeziboii):

repost on a new question board please

OpenStudy (anonymous):

dont you curse watch that mouth of yours

OpenStudy (mathmale):

Looks as though you don't understand that I was trying to help you by asking you questions. "you should know math-male" If you'd like to learn how to handle this material, why not follow the suggestions of others who are trying to help you?

OpenStudy (anonymous):

yeah but you should know

OpenStudy (anonymous):

you are a math male supposely

OpenStudy (anonymous):

Asiyah

OpenStudy (anonymous):

repost the question ina new board